# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 3.1 Extension to side information
Subsequent approaches were developed to take advantage of scenarios with available side information. Roughly speaking, these methods use side information to weight or rescore the hypotheses, placing higher weights or scores on hypotheses that are believed to be false nulls. Such procedures can be either ‘predetermined’, that is, the weighting or scoring is determined before the actual testing procedure takes place, or ‘adaptive’, in which the weighting or scoring is selected according to the data at hand and therefore often leads to greater statistical power.
@@
AdaFDR (2019) Asymptotic FDP Fast
Table 1: Summary of some popular and recent p-value based multiple testing procedures that use side information, indicating each method’s type-1 error control and speed.
@@
Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives [33, 8], Z-value Adaptive Procedures (ZAP) [35] and AdaFDR [51]. Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed. Notably, none of these mentioned tools offers a fast finite type-1 error control, and as we will see, some of these methods have variable power when considering a range of simulated and real datasets. We next provide some background on each of those methods.
AdaFDR splits the pairs $(p_{i},\mathbf{x}_{i})_{i=1}^{m}$ into two folds. For each fold, the method fits a mixture model to optimally rescore the hypotheses, and the model is then applied to the other fold where a list of discoveries is obtained. The two list of discoveries are joined together and then reported. AdaFDR’s mixture model is a composition of Gaussian components to model local ‘bumps’ of genuine signals (false nulls) within the data plus an additional component that captures the monotonic relationship between the side information and the signals.
AdaPT iteratively prunes a candidate set of hypotheses $\mathcal{S}$, similar to Adaptive Knockoffs, which was derived from it. At each iteration, AdaPT fits a two-group mixture model on a strict set of information: the p-values outside $\mathcal{S}$, the masked p-values in $\mathcal{S}$ given by $\{p_{i}\wedge(1-p_{i}):i\in\mathcal{S}\}$ where $1-p_{i}$ is referred to as the mirrored p-value, the side information $\mathbf{x}$, as well as #$\#\{p_{i}>1/2:i\in\mathcal{S}\}$ and #$\#\{p_{i}<1/2:i\in\mathcal{S}\}$. It then removes the hypothesis from $\mathcal{S}$ that is deemed most likely to be a true null based on the fitted model. To fit the mixture model, the EM algorithm is used. The M-step of the EM algorithm results in a regression problem which allows the user to flexibly use the regression method of their choice. The authors considered a generalized additive model (GAM), a generalized linear model (GLM), and a generalized linear model with $L_{1}$ regularization (GLMnet). If we identify each p-value $p_{i}>1/2$ with a label $L_{i}=-1$ and each p-value $p_{i}<1/2$ with a label $L_{i}=1$, then AdaPT terminates this pruning procedure when the estimated FDR, as in Equation (4), is $\leq\alpha$. The rationale is that for true null hypotheses, the p-values are usually uniformly distributed, so #$\#\{L_{i}=-1:i\in\mathcal{S}\}$ estimates the number of true null hypotheses with positive labels, #$\#\{L_{i}=1:i\in\mathcal{S}\}$. It then reports the remaining positively labeled hypotheses in $\mathcal{S}$ [^3] AdaPT and AdaPTg do not use labels explicitly — we introduce those here for notational convenience.
AdaPTg generalizes AdaPT by constructing asymmetric regions to define the positive and negative labels. For example, the p-values $p_{i}\in\left(0.3,0.9\right)$ can be identified with a label $L_{i}=-1$ and the p-values $p_{i}<0.3$ can be identified with a label $L_{i}=1$. One advantage of defining these asymmetric regions, is to avoid overestimating the number of true null hypotheses in cases where their p-values are not uniformly distributed and there is a concentration of true null p-values near 1. Subsequently, the estimated FDR needs to be adjusted to account for the fact that we expect twice as many true null p-values with negative labels than positive labels. Moreover, the masked p-values are also no longer $p_{i}\wedge(1-p_{i})$, since $1-p_{i}$ no longer ‘correctly mirrors’ the values in $\left(0.3,0.9\right)$ onto $(0,0.3)$ and therefore needs adjustment. In essence, the rest of the procedure remains the same. Lastly, AdaPT-GMMg replaces the two group mixture model in AdaPT/AdaPTg for a Gaussian mixture model (GMM). Similar to AdaPT/AdaPTg, it also allows the user to choose from a collection of regression methods to fit the model at the M-step of the EM algorithm.
Lastly, ZAP-asymp directly fits a beta mixture model to the data but is only able to offer asymptotic FDR control. Importantly, ZAP-asymp can operate directly on the test statistic instead of just the corresponding p-values to rerank the hypotheses. This is particularly advantageous for two-sided tests because the test statistic is mapped to its p-value in a non-bijective manner, losing information regarding the sign of the original test statistic along the way. Note that AdaPT-GMMg can also operate on test statistics, but in this paper, we focus on the p-value information.
@@
The list of methods we examine here is incomplete. Others include Independent Hypothesis Weighting [25, 24] which splits the p-values and side information into several folds and learns a weight for each hypothesis within that fold using out-of-fold information. This was shown to control the FDR asymptotically, and with some later adjustments to the method, finite sample FDR control was achieved. Another is Structure Adaptive Benjamini Hochberg Algorithm (SABHA) [37], which performs a reweighting of the p-values (similar to IHW) with finite sample FDR control. In addition, ZAP-finite is the analogue of ZAP-asymp which offers finite-sample FDR control. It uses the same principles of the AdaPT methods where each hypothesis is pruned one at a time. However, in previous simulations and real data experiments, IHW and SABHA appeared to be consistently outperformed by some of the methods in Table 1. Because we consider the same simulations and real data experiments, we did not include those methods in our comparisons. In addition, ZAP-finite is presented as a less powerful alternative to ZAP-asymp with the computational costs of the AdaPT methods. Given our comparisons to ZAP-asymp and the AdaPT methods, we decided to exclude ZAP-finite from our comparisons.
4 RESET (REScoring via Estimating and Training)
@@
RESET is designed to control the FDR or the FDP in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm S4. In Section S.3, we prove that RESET controls the FDR/FDP in the finite-sample setting. For simplicity, we use the target-decoy terminology, i.e., we refer to a hypothesis with $L_{i}=1$ as a ‘target win’ or a target for short, and $L_{i}=-1$ as a ‘decoy win’ or a decoy for short.
4.1 RESET outline
@@
    Intrinsically, RESET’s input consists of the labels, winning scores, and side information $(L,W,\mathbf{x})$. Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
@@
    where $0<a\leq 1/2\leq b_{1}<b_{2}\leq 1$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF. Hypotheses with p-values outside of $[0,a)\cup(b_{1},b_{2}]$ are thrown out. The default is $a=b_{1}=1/2$ and $b_{2}=1$.
@@
    RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout). The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo-targets is the set of all estimating decoys and target wins. The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo-target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
@@
    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores, and side information $(\tilde{L},W,\mathbf{x})$. Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels indicate true null hypotheses.
    The output of this step is a rescoring of the training decoys and pseudo-targets, denoted as $\tilde{W}$. Ideally, $\tilde{W}$ scores many of the false nulls among the pseudo-targets higher than they were scored originally. We provide a specific framework that we developed for this crucial step in Section 4.3.
@@
    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo-targets are revealed.
@@
    RESET then determines its list of discoveries by applying to the pseudo-targets using the original labels $L$ and the new scores $\tilde{W}$ either
    • SSS+ (Algorithm S1) at the desired level $\alpha$ for FDR control, or
@@
    • FDP-SD (Algorithm S2) at the desired level $\alpha$ and confidence parameter $\gamma$ for FDP control.
@@
    The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$. This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$. Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1; however, RESET uses approximately half of the decoys for training purposes (by default). Since those decoys are subsequently thrown out, this parameter needs to be adjusted. Indeed, assuming that the labels $L$ of the true nulls are i.i.d. uniform ±$\pm 1$ RVs, and $s=1/2$, we set the parameter $c=2/3$, so that the expected target-decoy ratio is 2. More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=\frac{c_{0}}{1-s\cdot(1-c_{0})}$. For example in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$. This choice is made rigorous in Lemma S1.
@@
We note that the above application of $\Phi^{-1}$ in Equation (5) is not strictly necessary for type-1 error rate control. We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$ because we found that doing so makes the semi-supervised learning task of Step 3 to be generally less variable and more powerful (data not shown).
@@
Crucially in Step 4, the training decoys are thrown out. This is important since during the training phase of Step 3, the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo-targets. Hence, true null targets will tend to score higher than the training decoys making the latter unsuitable for estimating the number of true nulls among the top scoring hypotheses.
4.2 RESET controls the FDR or FDP
To establish RESET’s control of the FDR or FDP in both the competition and p-value settings, we generalize Assumption 2 when the distribution of the true null labels $\{L_{i}:i\in N\}$ is non-uniform. This is given below:
@@ 4.2 RESET controls the FDR or FDP
Let $N$ be the indices of the true null hypotheses. The labels $\{L_{i}:i\in N\}$ are i.i.d. ±$\pm 1$ random variables with $\mathbb{P}(L_{i}=1)\leq c_{0}$ and are independent of all the scores $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses.
In the p-value setting, Assumption 3 is satisfied under some mild conditions regarding the distribution of the true null p-values. The value of $c_{0}$ is determined by the cutoff regions $[0,a)$ and $(b_{1},b_{2}]$ which are used to assign the positive or negative labels (Equation (5)). This is given by the following Lemma originally established by Chao and Fithian [8].
Lemma 1 (Chao and Fithian).
Assume that the true null p-values have a non-decreasing density and that they are mutually independent and that they are independent of the false null p-values as well as of all the side information. Then Assumption 3 is satisfied with $c_{0}=\frac{a}{a+b_{2}-b_{1}}$.
The above lemma is phrased differently by Chao and Fithian and is presented as Lemma A.1 in their paper, but it is materially the same. Most of the time, we will consider symmetric cutoff regions $[0,1/2)$ and $(1/2,1]$. For completeness, we provide a proof of this lemma in Supplementary Section S.2 using our notation.
In practice, Assumption 3 (or the conditions of Lemma 1) needs to be verified. We provide reasonable heuristic justification of this assumption in our real data applications in Sections 5.2 and 6.2. We are now ready to state our two main theorems.
@@
Under Assumption 3 RESET controls the FDP at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
We defer the proof of RESET’s FDR and FDP control to Supplementary Section S.3. The proof requires an intermediate lemma (Lemma S1), which states that the labels $L$ of the true null pseudo-targets in Step 4 remain i.i.d. ±$\pm 1$ RVs independently of everything else, even after sampling the training decoys in Step 2 and the application of the machine learning model in Step 3.

### v3

@@ 3.1 Extension to side information
Table 1: Summary of some popular recent p-value based multiple testing procedures that use side information.
@@
AdaFDR (2019) Asymptotic FDX Fast
@@
More recent works harness side information in a data-adaptive way to further boost statistical power. Such adaptive methods include Adaptive P-value Thresholding (AdaPT) and its derivatives (Lei and Fithian (2018); Chao and Fithian (2021)), Z-value Adaptive Procedures (ZAP) (Leung and Sun (2022)) and AdaFDR (Zhang et al. (2019)). Table 1 provides a qualitative summary of each method in terms of its type-1 error control, its error control guarantees, and its computational speed. ZAP-asymp and AdaPT-GMMg also have the advantage of operating directly on the test statistic instead of the corresponding p-values, which can boost power. In this paper, we will focus our comparisons based on the p-value information. Notably, none of these mentioned tools offer a fast finite type-1 error control (FDR or FDX), and as we will see, some of these methods have variable power when considering a range of simulated and real datasets. We provide a description of these methods in Section C.1.
@@
The list of methods we examine here is incomplete. Others include Independent Hypothesis Weighting (Ignatiadis et al. (2016); Ignatiadis and Huber (2021)) and Structure Adaptive Benjamini Hochberg Algorithm (SABHA) (Li and Barber (2019)) which both adaptively weight the p-values using side information. In addition, ZAP-finite is the analogue of ZAP-asymp which offers finite sample FDR control. It uses the same principles of the AdaPT methods where each hypothesis is pruned one at a time. However, in previous simulations and real data experiments, IHW and SABHA appeared to be consistently outperformed by some of the methods in Table 1 of the main text. Because we consider the same simulations and real data experiments, we did not include those methods in our comparisons. In addition, ZAP-finite is presented as a less powerful alternative to ZAP-asymp with the computational costs of the AdaPT methods. Given our comparisons to ZAP-asymp and the AdaPT methods, we decided to exclude ZAP-finite from our comparisons.
4 Methods
@@
RESET is designed to control the FDR or the FDX in either the competition or p-value based setting. We next describe RESET’s main steps, with further details given in Algorithm A4.
4.1 RESET
@@
    Intrinsically, RESET’s input consists of the labels, $L=(L_{i})_{i=1}^{m}$, the winning scores, $W=(W_{i})_{i=1}^{m}$, and side information, $\mathbf{x}=(\mathbf{x}_{i})_{i=1}^{m}$. Hence, in the p-value setting, each $p_{i}$ is first converted to a pair $(L_{i},W_{i})$ in the following way:
@@
    where $0<a\leq b_{1}<b_{2}\leq 1$ and $a\leq 1/2$ determine the cutoff regions for defining a positive and negative label, and $\Phi$ is the standard normal CDF (the motivation for the introduction of $\Phi$ is discussed below). Hypotheses with p-values outside of $0,a)\cup(b_{1},b_{2}]$ are thrown out. The default is $a=b_{1}=1/2$ and $b_{2}=1$.
@@
    RESET independently and randomly assigns each winning decoy as a training decoy with probability $s$ (we used $s=1/2$ throughout). The complementary set of decoy wins defines the estimating decoy set, and the set of pseudo targets is the set of all estimating decoys and target wins. The pseudo labels, $\tilde{L}_{i}$, denotes whether the $i$th hypothesis is a pseudo target ($\tilde{L}_{i}=1$) or a training decoy ($\tilde{L}_{i}=-1$).
@@
    Next, RESET applies a user-selected semi-supervised machine learning model which uses the pseudo labels, winning scores, and side information $(\tilde{L},W,\mathbf{x})$. Note that this is a semi-supervised task because the positive pseudo labels typically contain a mixture of true and false nulls, while the negative pseudo labels are mostly comprised of null hypotheses.
    The output of this step is a rescoring of the training decoys and pseudo targets, denoted as $\widetilde{W}$. Ideally, $\widetilde{W}$ scores many of the false nulls among the pseudo targets higher than they were scored originally. We provide a specific framework that we developed for this crucial step in Section 4.4.
@@
    With the training complete, the training decoys are thrown out, and the original labels $L$ of the remaining pseudo targets, $J:=\{i\in[m]:\tilde{L}_{i}=1\}$, are revealed.
@@
    Using the original labels $(L_{i})_{i\in J}$ and the new scores $(\widetilde{W}_{i})_{i\in J}$ RESET then determines its list of discoveries by applying either
    1. (a)
      SSS+ (Algorithm A1) at the desired level $\alpha$ for FDR control, or
    2. (b)
@@
      FDP-SD (Algorithm A2) at the desired level $\alpha$ and confidence parameter $\gamma$ for FDX control.
@@
    The two algorithms, SSS+ and FDP-SD, require an additional parameter $c$. This parameter is used to define the expected ratio of the number of null targets to decoys, $\frac{c}{1-c}$. Typically, this is set to $c=1/2$ for an expected target-decoy ratio of 1; however, RESET uses approximately half of the decoys for training purposes (by default). Since those decoys are subsequently thrown out, this parameter needs to be adjusted. Indeed, assuming that the labels $L$ of the true nulls are i.i.d. uniform $\pm 1$ RVs, and $s=1/2$, we set the parameter to $c=c_{e}:=2/3$, so that the expected target-decoy ratio is $\frac{c}{1-c}=2$. More generally, if we know that for a true null $\mathbb{P}(L_{i}=1)\leq c_{0}$, then we set $c=c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})}$. For example, in the p-value setting, $c_{0}=\frac{a}{a+b_{2}-b_{1}}$. This choice is made rigorous in Lemma 1.
@@
We note that the above application of $\Phi^{-1}$ in Equation (4) is not strictly necessary for type-1 error rate control. We apply $\Phi^{-1}$ to the p-value $p_{i}$ or the mirrored p-value, $(b_{2}-p_{i})\cdot\frac{a}{b_{2}-b_{1}}$, because we found that doing so makes the semi-supervised learning task of Step (3) to be generally less variable and more powerful (data not shown).
@@
Crucially, in Step (5) the training decoys are thrown out. This is important since during the training phase of Step (3), the semi-supervised machine learning model will have presumably learned to discriminate between the training decoys and the pseudo targets. Hence, true null targets will tend to score higher than the training decoys making the latter unsuitable for estimating the number of true nulls among the top scoring hypotheses.
4.2 Other related work and contributions
Recently, Marandon et al. (2024) developed a method called AdaDetect in the application of novelty detection. Given a sample of null observations $(Y_{1},\dots,Y_{n})$ with common marginal distribution $P_{0}$, the objective is to detect which observations in a test set $(X_{1},\dots,X_{m})$ are drawn from a different marginal distribution, i.e., $X_{i}\not\sim P_{0}$. AdaDetect uses a similar strategy to that of RESET by splitting the null observations into two parts, $(Y_{1},\dots,Y_{k})$ and $(Y_{k+1},\dots,Y_{n})$. The first part is used to assist in ranking the joint observations $(Y_{k+1},\dots,Y_{n},X_{1},\dots,X_{m})$ while the second part is used to construct empirical p-values (also called conformal p-values (Bates et al. (2023))) for each test set observation. Then, the BH procedure (or one of its $\pi_{0}$ derivatives) is applied.
The critical assumption that guarantees AdaDetect’s FDR control is that the null observations $(Y_{1},\dots,Y_{n})$ and the null test set observations $(X_{i}:X_{i}\sim P_{0})$ are exchangeable, conditioned on $(X_{i}:X_{i}\not\sim P_{0})$. A natural question is whether a similar set of assumptions can be applied to the multiple hypothesis setting, so that an AdaDetect-like approach can be used to control the FDR. The answer is no, and we show in Section F.1 that applying AdaDetect’s idea can break FDR control.
We recently developed Percolator-RESET (Freestone et al. (2025)) that combined the RESET approach with Percolator, a popular semi-supervised algorithm designed specifically for FDR control in peptide detection from mass spectrometry data (Käll et al. (2007)). This manuscript includes several novel contributions over Freestone et al. (2025). First, we introduce the general RESET algorithm, which is applicable to both p-value and competition-based multiple testing scenarios. Second, we also consider FDX control and provide theoretical justification of RESET’s finite sample FDR or FDX control (thereby also providing the previously-missing theoretical proof of Percolator-RESET’s FDR control). Third, we introduce “RESET Ensemble” as the new semi-supervised engine of RESET’s Step (c). Fourth, we conduct extensive simulated and real data experiments comparing RESET Ensemble with competing methods (including a comparison of the new RESET Ensemble to Adaptive Knockoffs in the peptide detection context). Fifth, we explore FDX control with side information for peptide detection.
4.3 RESET controls the FDR or FDX
To establish RESET’s control of the FDR or FDX in both the competition and p-value settings, we relax Assumption 2 so that it can be applied to the case where the true null labels are no longer necessarily uniformly or even identically distributed:
@@ 4.3 RESET controls the FDR or FDX
Let $N$ be the indices of the true null hypotheses. Conditioned on $W$, the side information $\mathbf{x}$, and the labels of the false null hypotheses, the true null labels $(L_{i})_{i\in N}$ are independent $\pm 1$ random variables with $\mathbb{P}(L_{i}=1)=:c_{i}\leq c_{0}$.
We discuss the applicability of this assumption in Sections 5 and 6 where we apply RESET in both the competition and p-value settings.
RESET’s FDR/FDX control hinges on the following Lemma, which essentially states that the labels $L$ of the true null pseudo targets in Step (5) are conditionally independent $\pm 1$ RVs—even after we sample the training decoys in Step (2) and the application of the machine learning model in Step (3).
Lemma 1.
Let $\mathcal{G}:=\sigma\left(W,\mathbf{x},(L_{i})_{i\not\in N},(\tilde{L}_{i})_{i\in[m]}\right)$, $N$ be the indices of the true nulls, and $J:=\{i\in[m]:\tilde{L}_{i}=1\}$. Then, under Assumption 3, conditioned on $\mathcal{G}$, $(L_{i})_{i\in J\cap N}$ are independent $\pm 1$ RVs with
$\mathbb{P}(L_{i}=-1\mid\mathcal{G})\geq 1-c_{e},$
where $c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})}$.
Proof.
Note that we can express $\tilde{L}_{i}$ as $\tilde{L}_{i}=f(L_{i},B_{i})$ where $f$ is a deterministic function and $B_{i}$ are Bernoulli(s)(s) RVs that are independent given everything else. It follows from $\tilde{L}_{i}=f(L_{i},B_{i})$ and Assumption 3 that, conditioned on $\mathcal{G}_{0}:=\sigma\left(W,\mathbf{x},(L_{i})_{i\not\in N}\right)$, the pairs in $\{(L_{i},\tilde{L}_{i}):i\in N\}$ are independent of one another.
Moreover, from $\tilde{L}_{i}=f(L_{i},B_{i})$ and $(L_{i})_{i\not\in N}\in\mathcal{G}_{0}$ we find that these pairs are still independent conditioned on $\sigma\left(\mathcal{G}_{0},(\tilde{L}_{i})_{i\not\in N}\right)$. This independence in turn implies that $(L_{i})_{i\in N}$ are independent if we further condition on all the pseudo labels, i.e., on $\mathcal{G}=\sigma\left(\mathcal{G}_{0},(\tilde{L}_{i})_{i\in[m]}\right)$. Since $J$ is fixed conditioned on $\mathcal{G}$, the labels $(L_{i})_{i\in J\cap N}$ are independent conditioned on $\mathcal{G}$, establishing the first part.
It follows from Assumption 3 and the above discussion, that with $\mathcal{G}_{-i}:=\sigma\left(\mathcal{G}_{0},(\tilde{L}_{j})_{j\neq i}\right)$, $\mathbb{P}(L_{i}=-1\mid\mathcal{G}_{-i})\geq 1-c_{0}$. Consider $i\in N$ and an atom of $\mathcal{G}_{-i}$ where $W$, $\mathbf{x}$, $(L_{j})_{j\not\in N}$ and $(\tilde{L}_{j})_{j\neq i}$ are all fixed. Then construing the following as conditional on that atom we have
$\geq(1-c_{0})\cdot(1-s).$
Hence, under the same assumptions
$=\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot\mathbb{P}(L_{i}=-1)}$
$\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}$
$=1-c_{e}.$
Additionally, clearly
$\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=-1)=1\geq 1-c_{e}.$
Therefore, under the same assumptions, the following holds a.s.
$\mathbb{P}(L_{i}=-1\mid\mathcal{G})=\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i},\mathcal{G}_{-i})\geq 1-c_{e}.$
It is worth noting that Lemma 1 is not true for an arbitrary splitting of the decoys into the training and estimating sets, e.g., splitting the decoys into two halves of equal size. To see this, suppose that all hypotheses are true nulls, i.e., $N=[m]$. Then by observing the number of training decoys, we know that the number of true null decoys in $J$ must be the same, and the labels $(L_{i})_{i\in J\cap N}$ are no longer independent. We are now ready to state our two main theorems.
@@
Under Assumption 3 RESET controls the FDX at the user-specified threshold $\alpha$ and confidence $1-\gamma$.
The proofs of Theorem 1 and 2 follow from Lemma 1 and rely on a technical discussion when the labels $(L_{i})_{i\in J\cap N}$, conditioned on $\mathcal{G}$, are non-identical, which we give in Section B.1. To briefly see the connection of Lemma 1 to Theorems 1 and 2, note that the labels $(L_{i})_{i\in J\cap N}$ are independent $\pm 1$ RVs even after observing the machine learning scores $\widetilde{W}$ in Step (3) and the false null labels $(L_{i})_{i\notin N}$. If the labels are identical, i.e., $\mathbb{P}(L_{i}=-1\mid\mathcal{G})=1-c_{e}$, then the assumptions of SSS+ in Barber and Candès (2015) and FDP-SD in Luo et al. (2023) are satisfied, and RESET controls the FDR or FDX, respectively. Intuitively, in the non-identical case, i.e., $\mathbb{P}(L_{i}=-1\mid\mathcal{G})\geq 1-c_{e}$, FDR or FDX control should be more conservative using these procedures.